### v2

# Learning Graph Quantized Tokenizers

Abstract
Transformers serve as the backbone architectures of Foundational Models, where domain-specific tokenizers allow them to adapt to various domains. Graph Transformers (GTs) have recently emerged as leading models in geometric deep learning, outperforming Graph Neural Networks (GNNs) in various graph learning tasks. However, the development of tokenizers for graphs has lagged behind other modalities. To address this, we introduce GQT (Graph Quantized Tokenizer), which decouples tokenizer training from Transformer training by leveraging multi-task graph self-supervised learning, yielding robust and generalizable graph tokens. Furthermore, the GQT utilizes Residual Vector Quantization (RVQ) to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities. By combining the GQT with token modulation, a Transformer encoder achieves state-of-the-art performance on 20 out of 22 benchmarks, including large-scale homophilic and heterophilic datasets.
1 Introduction
Following the success of Transformers (Vaswani et al., 2017) in Natural Language Processing (Devlin et al., 2019; Brown et al., 2020) and Computer Vision (Dosovitskiy et al., 2021), Graph Transformers (GTs) (Dwivedi and Bresson, 2020; Ying et al., 2021; Rampášek et al., 2022; Shirzad et al., 2023; Chen et al., 2023; Wu et al., 2022) have emerged as strong models in geometric deep learning. In contrast to message-passing Graph Neural Networks (GNNs), which are inherently constrained by strong locality inductive biases (Battaglia et al., 2018; Veličković et al., 2018; Hou et al., 2020; Hamilton et al., 2017; Kipf and Welling, 2017; Zheng et al., 2024), GTs exhibit greater expressivity due to their capacity to capture long-range interactions between nodes (Ma et al., 2023; Kim et al., 2022; Zopf, 2022). This is particularly beneficial in heterophilic settings where local alignment does not hold (Fu et al., 2024). This dichotomy highlights a fundamental trade-off between GNNs, which focus on local neighborhood aggregation, and GTs, which employ pairwise attention to model global graph structures. A natural question arises: can we synergistically integrate the strengths of both approaches to leverage the complementary benefits of local and global representations? Specifically, is it possible to harness the locality-aware representations learned by GNNs to construct discrete tokens, thereby enabling GTs to operate efficiently while still capturing salient graph properties?
GTs require consideration of both graph structure and features, as nodes with identical features will otherwise be projected into the same representation regardless of their surrounding structures (Hoang et al., 2024). There are three general approaches to address this limitation (Hoang et al., 2024): (1) node feature modulation, which involves injecting structural information into the node features; (2) context node sampling, where a sampling strategy is used to construct a sequence over the neighbor nodes; and (3) modifying the architecture of a vanilla Transformer to directly incorporate structural biases. Given that Transformers are universal approximators of sequence-to-sequence functions (Yun et al., 2020) and considering the rapid developments in efficient implementation of Multi-Head Attention (MHA) module (Dao et al., 2022; Liu et al., 2024), which enables longer context sizes (Reid et al., 2024), we argue that a well-designed graph tokenizer can allow a vanilla Transformer to efficiently process even large-scale graphs. Recent studies on applying Large Language Models (LLMs) to Text-Attributed Graphs (TAGs) have shown surprisingly strong performance gains surpassing those of GNNs, suggesting that vanilla Transformers are indeed capable of effectively learning graph structures (Ye et al., 2024; Xu et al., 2025). Nonetheless, LLMs are not efficient at inference time. Our goal is to devise a lightweight and efficient graph tokenizater that allows vanilla Transformer encoders to learn graph structures effectively.
Tokenizers typically employ self-supervised objectives to abstract data into a sequence of discrete tokens, allowing Transformers to learn representations across various modalities as a unified stream of data. The discretization is usually achieved through vector quantization techniques (Van Den Oord et al., 2017; Lee et al., 2022), which offer several benefits, including: (1) significantly reduced memory requirements, (2) improved inference efficiency, (3) allowing Transformers to focus on long-range dependencies rather than local information, and (4) the capacity to learn more high-level representations due to a compact latent space (Yuan et al., 2021; Yu et al., 2022a). These advantages are particularly important in auto-regressive generative modeling, where quantized tokens allow Transformers to generate high-quality outputs in multiple modalities (Dubey et al., 2024; Lee et al., 2022; Dhariwal et al., 2020; Ramesh et al., 2021; Team, 2024). Despite its importance in other domains, tokenization remains under-explored for graph-structured data. To address this limitation, we propose the Graph Quantized Tokenizer (GQT), a novel approach that learns a hierarchical sequence of tokens over graphs using self-supervised objectives tailored to graph-structured data. More specifically, our contributions are as follows:
  • We propose a graph tokenizer that uses multi-task graph self-supervised learning to train a graph encoder, enabling it to fully capture local interactions and allowing the Transformer to focus on long-range dependencies.
  • Our approach adapts Residual Vector Quantization (RVQ) within the graph tokenizer to learn hierarchical discrete tokens, resulting in significantly reduced memory requirements and improved generalization capabilities.
  • We introduce a novel combination of semantic edges and random walks to facilitate access to long-range interactions, and employ hierarchical encoding and gating mechanisms to modulate the tokens and provide informative representations to the Transformer.
  • Through extensive experiments on both homophilic and heterophilic datasets, including large-scale and long-range benchmarks, we demonstrate that our tokenizer enables Transformer encoders to achieve state-of-the-art performance on 20 out of 22 benchmarks while substantially reducing the memory footprint of the embeddings.
2 Related Works
Graph Transformers (GTs) have shown promising performance on various graph learning tasks, surpassing GNNs on many benchmarks. Designing GTs can be broadly categorized into two directions (Hoang et al., 2024; Müller et al., 2024): (1) modifying the vanilla Transformer architecture to incorporate structural inductive biases, or (2) encoding the input graph to make it compatible with the vanilla Transformer. Early examples of the first approach include Graph Attention Network (GAT) (Veličković et al., 2018), which uses an attention module to compute pairwise node attention and masks the attention matrix based on connectivity information. Subsequent works have replaced the scaled-dot attention module with various structure-aware sparse attention modules (Rampášek et al., 2022; Bo et al., 2023; Ying et al., 2021; Deng et al., 2024; Wu et al., 2023b; Liu et al., 2023a; Chen et al., 2022; Dwivedi and Bresson, 2020; Shirzad et al., 2023; Ma et al., 2023). Graph Memory Network (GMN) (Khasahmadi et al., 2020) is an example of the second approach, which passes non-linear projections of node features and structural encoding to a Transformer-like model. Structural encodings such as Laplacian eigenvectors or Random walk-based encoding (Dwivedi et al., 2022a; Ma et al., 2023; Cantürk et al., 2024), allow injecting structural information directly into the node features. Some works use GNNs to encode local structure along with node features into embeddings that are passed to vanilla Transformers to capture long-range dependencies (Rong et al., 2020; Wu et al., 2021; Chen et al., 2023, 2022). Recent studies leverage LLMs, where graphs are represented through natural language, and an LLM performs graph-related tasks through in-context learning, instruction-tuning, or soft-prompting (Fatemi et al., 2024; Ye et al., 2024; He et al., 2024). For a detailed survey on GTs, see (Müller et al., 2024; Hoang et al., 2024).
Graph Tokenization provides GTs with rich node tokens that encapsulate both structural and semantic information. TokenGT (Kim et al., 2022) treats nodes and edges as independent tokens defined by their features, type identifiers, and structural encodings. NAGphormer (Chen et al., 2023) represents each node with $L$ tokens, where the $l^{th}$ token is the representation of the node from the $l^{th}$ hop aggregation. GraphiT (Mialon et al., 2021) defines a node token as the concatenation of its feature and representation from a graph convolutional kernel network (GCKN). VCR-Graphormer (Fu et al., 2024) expands the notion of node tokens to include sequences comprising the node feature and features of semantically and community-related neighboring nodes. SGT (Liu et al., 2023c) is a non-parametric tokenizer designed for molecular tasks, which unlike motif-based tokenizers (Zhang et al., 2021; Jin et al., 2018) or GNN pre-training methods (Xia et al., 2023), simplifies the tokenization process to a non-parametric graph operator without non-linearity. NodePiece (Galkin et al., 2022) is a knowledge-graph tokenizer that represents a target node as a hash of its top-k closest anchors, their distances, and relational context. While Vector Quantization (VQ) has been explored in other modalities (Van Den Oord et al., 2017; Lee et al., 2022; Yu et al., 2022b; Van Kempen et al., 2024; Li et al., 2024), its application in graph learning is limited. Notable exceptions include VQ-GNN (Ding et al., 2021), which uses quantized representations combined with a low-rank graph convolution matrix to avoid neighbor explosion problem, VQGraph (Yang et al., 2024), which employs VQ for distilling a GNN into an MLP, and NID (Luo et al., 2024a), which uses VQ to learn discrete node IDs for downstream prediction tasks.
3 Preliminaries
Messag-Passing GNNs. Let $\mathcal{G}$ denote the space of graphs. A graph $g\in\mathcal{G}$ is defined as $\left(\mathcal{V},\mathcal{E},\textbf{X},\textbf{E}\right)$ where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. $\textbf{X}\in\mathbb{R}^{\left|\mathcal{V}\right|\times d_{x}}$ represents the node features of dimension $d_{x}$, and $\textbf{E}\in\mathbb{R}^{\left|\mathcal{V}\right|\times\left|\mathcal{V}\right|\times d_{e}}$ represents the edge features of dimension $d_{e}$. A message-passing GNN takes $g$ as input and learns representations $h^{l}_{v}$ for $v\in\mathcal{V}$ ($h^{0}_{v}=x_{v}$) in each layer $l$ as follows (Gilmer et al., 2017):
$h^{l}_{v}=f^{l}_{\theta}\left(h^{l-1}_{v},g^{l}_{\phi}\left(\left\{\left(h^{l-1}_{v},h^{l-1}_{u},e_{uv}\right)|u\in\mathcal{N}_{v}\right\}\right)\right)$ (1)
where $f_{\theta}$ and $g_{\phi}$ are known as combine and aggregate functions, respectively. $\mathcal{N}_{v}$ denotes the set of immediate neighbors of the node $v$. Once the node representations are computed, we can perform various tasks including node classification as $\text{MLP}\left(h_{v}\right)$, edge prediction as $\text{MLP}\left(h_{u}\odot h_{v}\right)$, or graph classification as $\text{MLP}\left(\mathcal{R}\left(\left\{h_{u}|u\in\mathcal{V}\right\}\right)\right)$, where $\mathcal{R}$ is a pooling (readout) function.
Graph Transformers use a tokenizer $T_{v}=\mathcal{T}_{\psi}\left(\mathcal{N}(v)\right)$ to map each node $v\in\mathcal{V}$ into a sequence of tokens $T_{v}$ by considering a notion of neighborhood $\mathcal{N}$. The simplest design is when $\mathcal{N}$ is zero-hop neighborhood (i.e., the node itself) and $\mathcal{T}_{\psi}$ is a node feature lookup function. The neighborhood $\mathcal{N}$ can be extended to include the node’s ego network (Zhao et al., 2021) or top-k Random Walk based neighbors (Fu et al., 2024), and $\mathcal{T}_{\psi}$ can be enhanced to representations from a GNN (Chen et al., 2023). Node tokens along with positional encodings (PE) are passed to the Transformer as $h^{0}_{v}=\left[T_{v}||\text{PE}\left(v\right)\right]$. The representations in the $l^{th}$ layer of a Transformer encoder are computed as:
$h^{l}_{v}$ $=\text{LN}\left(\text{MHA}\left(\text{LN}\left(h^{l-1}_{v}\right)\right)+h^{l-1}_{v}\right)$ (2)
$h^{l}_{v}$ $=h^{l}_{v}+\text{MLP}\left(h^{l}_{v}\right)$ (3)
[…]
Vector Quantization projects embeddings $\textbf{X}\in\mathbb{R}^{n\times d_{x}}$ into a more compact space of codebooks $\textbf{C}\in\mathbb{R}^{k\times d_{c}}$, where $k\ll n$. The codebooks can be learned by minimizing various objectives such as K-means clustering. The new representation of $x_{i}$ is then computed as (Van Den Oord et al., 2017):
[…]
Building upon this concept, Residual-VQ (RVQ) (Lee et al., 2022) extends VQ to a sequence of codebooks, where each consecutive codebook quantizes the residual error from the previous codebook, i.e., $r_{i}=z_{i}-c_{k}$. This hierarchical approach constructs a multi-level quantized representation, enhancing the overall quantization quality. More details of RVQ are included in Appendix 9.
4 Self-Supervised Graph Tokenization
4.1 Tokenizer Properties
Our goal is to design a graph tokenizer that learns node tokens that exhibit three key characteristics:
Modeling Local Interactions. The tokens should encapsulate local interactions, allowing the Transformer to focus on long-range dependencies. This is analogous to Vision Transformers (ViTs), where the Transformer attends to image patches instead of pixels (Dosovitskiy et al., 2021; Liu et al., 2021). To achieve this, we leverage GNNs as the tokenizer encoder to model local interactions in the representation space (Battaglia et al., 2018). Our design accommodates various GNN layer choices without constraints; for simplicity, we opt for a GAT encoder (Veličković et al., 2018).
Memory Efficiency. The tokens also should be compact to facilitate efficient memory usage. To achieve this, we introduce a Residual-VQ (RVQ) (Lee et al., 2022) layer to quantize the GNN representations into a sequence of discrete tokens. Quantization not only helps with generalization due to its regularization effect but also significantly reduces memory usage. Using an RVQ with $c$ codebooks (typically $c=\{2,\cdots,8\}$), a graph with feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{x}}$ can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{N\times c}$ and codebook representation of $\textbf{C}\in\mathbb{R}^{c\times K\times d_{c}}$, where $c$ is the number of codebooks (i.e., levels of quantization), $K$ is the codebook size, and $d_{c}$ is the code dimension. To illustrate the benefits of this approach, consider a graph with $10^{6}$ nodes and a feature dimension of 1024 ($\textbf{X}\in\mathbb{R}^{10^{6}\times 1024}$). Using an RVQ with 3 codebooks and a codebook size of 256, this graph can be represented as $\textbf{X}_{Q}\in\mathbb{N}^{10^{6}\times 3}$ plus $\textbf{C}\in\mathbb{R}^{3\times 256\times 1024}$, resulting in a 270-fold reduction in memory.
Robustness and Generalization. The tokens should be robust and generalizable. To achieve this, we rely on graph self-supervised learning. Self-supervised representations have been shown to be more robust to class imbalance (Liu et al., 2022) and distribution shift (Shi et al., 2023), while also capturing better semantic information (Assran et al., 2023) compared to representations learned through supervised objectives. Moreover, self-supervised graph representations have demonstrated superior performance on downstream tasks compared to representations learned in a fully supervised manner, indicating better generalization capabilities (Hu et al., 2020b; Sun et al., 2020; You et al., 2020; Fu et al., 2020; You et al., 2021; Hassani and Khasahmadi, 2020; Veličković et al., 2019; Zhu et al., 2020b). Additionally, multi-task learning with self-supervised objectives has been shown to achieve better performance on downstream tasks (Doersch and Zisserman, 2017; Ghiasi et al., 2021). To leverage these benefits, we propose training the GNN encoder with three self-supervised objectives. Unlike RQ-VAE (Lee et al., 2022), which uses reconstruction as its primary objective, we employ graph-specific objectives to capture the nuances of both structure and features.
Figure 1: Overview of our proposed Graph Quantization Transformer (GQT) consisting of three main components: (1) a GNN to encode local interactions, (2) vector quantization for compact representation, and (3) generative and contrastive heads for robust representation learning. We also utilize a Transformers encoder to model long-range interactions. We augment the graph with semantic edges (dashed lines) and generate a sequence for each node based on Personalized PageRank scores. We then modulate the tokens through hierarchical encoding and structural gating, and feed them into the Transformer and aggregate the learned representations through an attention module before passing it to the classification head.
4.2 Training
To capture different aspects of information, we employ a multi-task learning framework that leverages three distinct families of graph self-supervised objectives: student-teacher distillation (Thakoor et al., 2022), masked autoencoding (Hou et al., 2022), and Infomax (Veličković et al., 2019). We also introduce a commitment loss (Van Den Oord et al., 2017) to enforce alignment between learned node representations and the codebook representations. Specifically, the GNN encoder is trained through gradient descent to minimize a loss function comprising of three terms, where $\beta$ is the loss weight:
$\mathcal{L}=\mathcal{L}_{\text{dgi}}+\mathcal{L}_{\text{gmae2}}+\beta\mathcal{L}_{\text{commit}}$ (5)
The first term is the Deep Graph Infomax (DGI) (Veličković et al., 2019) objective, which maximizes mutual information (MI) between node representations and graph (sub-graph) representations, based on the Jensen-Shannon divergence between the joint and product of marginals as follows:
$\mathcal{L}_{\text{dgi}}=\mathbb{E}\left(\sum\limits_{v\in g}\log\left(\mathcal{D}\left(h_{v},h_{g}\right)\right)+\sum\limits_{u\in\tilde{g}}\log\left(1-\mathcal{D}\left(h_{u},h_{g}\right)\right)\right)$ (6)
where $h_{u}$ is the representation of node $u$. $h_{g}$ is the global (sub-graph/graph) representation, computed as the mean of node representations. $\tilde{g}$ is the corrupted version of the original graph, with the same structure but randomly shuffled features, providing negative examples for contrastive learning. $\mathcal{D}\left(h_{u},h_{g}\right)=\sigma\left(h_{u}^{T}\textbf{W}h_{g}\right)$ is the discriminator that scores whether a node belongs to the graph, and is defined as a bilinear classifier.
The second term is the GraphMAE2 objective (Hou et al., 2023), which combines the generative loss of GraphMAE (Hou et al., 2022) with the teacher-(noisy)student distillation loss of BGRL (Thakoor et al., 2022). This combination enables the model to avoid overfitting and learn more semantic representations. The GraphMAE2 loss is computed as follows:
$\mathcal{L}_{\text{gmae2}}=\sum\limits_{v\in\tilde{g}}\left(1-\frac{x_{v}^{T}.\tilde{h}_{v}}{\|x_{v}^{T}\|.\|\tilde{h}_{v}\|}\right)^{\gamma}+\lambda\sum\limits_{v\in g}\left(1-\frac{h_{v}^{T}.\tilde{h}_{v}}{\|h_{v}^{T}\|.\|\tilde{h}_{v}\|}\right)^{\gamma}$ (7)
where $\tilde{g}$ is the masked graph, $\tilde{h}_{v}$ is the node representation of a masked node learned by the noisy student, $h_{v}$ is the corresponding node representation learned by the teacher over the original graph, and $\gamma\geq 1$ is a scaling factor. The teacher’s parameters are updated using an Exponential Moving Average (EMA) of the noisy student’s parameters.
The third term is the commitment loss, which encourages the representations to get close to their corresponding codebook embeddings within the RVQ layer. This loss is computed as:
$\mathcal{L}_{\text{commit}}=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in g}||h_{v}-\text{sg}\left[c_{k}\right]||_{2}$ (8)
where sg is the stop-gradient operator, and $c_{k}$ is the representation of the codebook that $h_{v}$ is assigned to (i.e., the centroid or prototype vector). Note that this loss only affects the node representations and does not update the codebooks.
To initialize and update the codebooks, we employ K-Means clustering and EMA with weight decay $\tau\in[0,1]$, respectively. Specifically, the codebooks are updated as follows:
$c_{k}^{t}=\tau c_{k}^{t-1}+(1-\tau)\frac{1}{|\mathcal{V}_{k}|}\sum\limits_{v\in\mathcal{V}_{k}}h_{v}$ (9)
where $\mathcal{V}_{k}$ is the set of nodes assigned to codebook $c_{k}$. This update rule allows the codebooks to adapt to the changing node representations while maintaining stability.
5 Graph Transformer
5.1 Graph Serialization
Once the tokenizer is trained, each node $v\in\mathcal{V}$ is mapped to $c$ discrete tokens: $T_{v}=\left[t^{v}_{1},\cdots,t^{v}_{c}\right]\in\mathbb{N}^{c}$ (i.e., $T_{v}=\textbf{X}_{Q}[v]$), encoding local interactions of that node. We then need to serialize the graph in order to input it to the Transformer.
Semantic Edges. To enable the Transformer to capture long-range interactions, the input should consist of a sequence of tokens from nodes that are likely to have long-range dependencies. To facilitate this, we first augment the graph with semantic edges denoted as $\mathcal{E}_{s}$, which are computed as follows:
[…]
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $x_{u}$ is the feature vector of node $u$, and $f$ is a projection function. We use cosine similarity as the similarity function and Principal Component Analysis (PCA) as the projection function. The semantic edge augmentation effectively creates sparse edges between each node and its k-nearest neighbors in the feature space, enhancing the model’s ability to recognize and utilize long-range dependencies.
Structural Serialization. We combine the semantic edges with the original edges and use Personalized PageRank (PPR) to generate a sequence per target node. This enriches the sequence with information beyond local interactions, allowing the Transformer to access potential long-range dependencies. We construct the sequence $S_{v}$ for each node $v$ as follows:
[…]
where $S_{v}=\left[t^{v}_{1}\cdots t^{v}_{c}\mid t^{u_{1}}_{1}\cdots t^{u_{1}}_{c}\mid\cdots\mid t^{u_{k}}_{1}\cdots t^{u_{k}}_{c}\right]$ is a sequence of length $c\times(k+1)$, comprising discrete tokens that represent the target node $v$, followed by discrete tokens of the top-$k$ relevant nodes to node $v$. These relevant nodes are determined based on PPR scores. Note that the computation of semantic edges and PPR sequences is performed only once as a pre-processing step, thereby reducing the computational overhead during training.
5.2 Token Modulation
Token Embeddings. There are $c\times K$ possible discrete tokens, where $c$ is the number of codebooks and $K$ is the codebook size. We randomly initialize an embedding matrix $\textbf{X}_{T}\in\mathbb{R}^{c\times K\times d_{x}}$, which is trained end-to-end alongside the Transformer. To further enrich the token representations, we introduce an additional token for each node that aggregating the embeddings of its assigned codebooks from the pretrained tokenizer:
$h_{c}^{v}=\sum\limits_{i=1}^{c}\textbf{C}[i,t_{i}^{v}]$ (12)
where $\textbf{C}[i,j]$ is the embedding corresponding to index j in the ith codebook. We found that adding this explicit aggregated token from the codebook leads to better performance compared to initializing $\textbf{X}_{T}$ directly with C. The input representation of the sequence for node $v$ is then defined as:
$S_{v}=\left[\textbf{X}_{T}[i,t^{v}_{i}]\operatorname*{\|}_{i=1}^{c}h_{c}^{v}\ \|\ \textbf{X}_{T}[i,t^{u_{1}}_{i}]\operatorname*{\|}_{i=1}^{c}\ h_{c}^{u_{1}}\ \|\cdots\|\textbf{X}_{T}[i,t^{u_{k}}_{i}]\operatorname*{\|}_{i=1}^{c}\ h_{c}^{u_{k}}\right]$ (13)
[…]
Structural Gating. In order to provide the Transformer with the global structural importance scores of the nodes within the sequence with respect to the target node, we introduce a gating mechanism over the input token embeddings as follows:
[…]
where we first apply a softmax function with temperature $\tau=1$ to normalize the PPR scores, and then multiply each node token’s representation by its corresponding normalized score.
Positional Encoding. We also introduce two trainable positional encodings to the input tokens. The first positional encoding enables the Transformer to distinguish between tokens from different nodes, while the second encoding, referred to as hierarchical encoding, allows the Transformer to recognize the hierarchy level of each token within the codebooks. We randomly initialize the positional encodings $\textbf{PE}\in\mathbb{R}^{(k+1)\times d_{x}}$ and $\textbf{HE}\in\mathbb{R}^{c\times d_{x}}$ and sum them with the encoding of their corresponding token. For example, the final encoding of the token $j$ of the node $i$ within the sequence is computed as: $x=\textbf{X}_{T}[j,t^{u_{i}}_{j}]+\textbf{PE}[i]+\textbf{HE}[j]$. Note that we did not use any structural encoding, such as Laplacian eigenvectors, as we did not observe any significant gains from them.
5.3 Transformer Encoder
We use $l$ layers of standard Transformer encoder with flash attention (Dao et al., 2022) to generate contextual representations per token in the sequence: $\textbf{H}^{(l)}\in\mathbb{R}^{(c+1)\times(k+1)\times d_{h}}$. We then aggregate the token representations for $j$-th node in the sequence by summing along the token dimension:
$\textbf{H}_{v_{j}}=\sum\limits_{i=1}^{c+1}\textbf{H}^{(l)}[i,j]\in\mathbb{R}^{(k+1)\times d_{h}}$ (15)
To obtain a single representation for the entire sequence, We further aggregate the representation using a linear attention layer:
$h=\sum\limits_{i=1}^{k+1}\alpha_{i}h_{i}\quad\text{where}\quad\alpha_{i}=\frac{\exp(\textbf{W}h_{i})}{\sum_{j}\exp(\textbf{W}h_{j})}$ (16)
We feed the resulting representation into a fully-connected classifier and train the model end-to-end using cross-entropy loss. Note that during inference, only the Transformer and classifier are utilized, as the tokenizer is pretrained and the sequences are pre-computed. Furthermore, since we only require discrete tokens and codebook embeddings, our approach enables efficient memory usage, regardless of graph size, allowing for efficient training and inference on large-scale graphs.
6 Experiments
We evaluate GQT on both medium- and large-scale graph learning tasks, encompassing 22 homophilic, heterophilic, and long-range benchmarks. We follow the established experimental protocols from previous works to ensure fair comparisons. Details of the datasets, experimental setup, and hyperparameters are provided in Appendices 10 and 11, respectively.
6.1 Comparison with State-of-the-Art
Long-Range Benchmarks. We use four datasets from the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022b), including the Peptides-Func dataset for graph classification with Average Precision (AP) metric, the Peptides-Struct dataset for graph regression with Mean Absolute Error (MAE) metric, the COCO-SP dataset for inductive node classification with macro F1 metric, and the PCQM-Contact for link prediction with Mean Reciprocal Rank (MRR) metric. We compare our results to baselines reported in (Wang et al., 2024). The results shown in Table 1 suggest that GQT is able to capture long-range dependencies and performs well on various graph prediction tasks.
Table 1: Mean performance on inductive long-range benchmarks over five runs.
Task Graph Classification Graph Regression Node Classification Link Prediction
Dataset Peptides-Func Peptides-Struct COCO-SP PCQM-Contact
#Graphs 15,535 15,535 123,286 529,434
Avg. #Nodes 150.94 150.94 476.88 30.14
Avg. #Edges 307.30 307.30 2,693.67 61.09
Metric AP ↑ MAE ↓ F1 ↑ MRR ↑
GCN 0.5930$\pm$0.0023 0.3496$\pm$0.0013 0.0841$\pm$0.0010 0.3234$\pm$0.0006
Exphormer 0.6258$\pm$0.0092 0.2512$\pm$0.0025 0.3430$\pm$0.0108 0.3587$\pm$0.0025
GPS 0.6535$\pm$0.0041 0.2500$\pm$0.0005 0.3412$\pm$0.0044 0.3337$\pm$0.0006
Graph-Mamba 0.6739$\pm$0.0087 0.2478$\pm$0.0016 0.3960$\pm$0.0175 0.3395$\pm$0.0013
GQT (Ours) 0.6903$\pm$0.0085 0.2452$\pm$0.0018 0.4007$\pm$0.0125 0.3427$\pm$0.0012
Table 2: Mean node classification accuracy on medium-scale homophilic datasets over five runs.
Dataset CoraFull CiteSeer PubMed Computer Photo CS Physics WikiCS
#Nodes 19,793 3,327 19,717 13,752 7,650 18,333 34,493 11,701
#Edges 126,842 4,522 88,651 491,722 238,163 163,788 495,924 216,123
#Features 8,710 3,703 500 767 745 6,805 8,415 300
#Classes 70 6 3 10 8 15 5 10
GNN GCN 61.76$\pm$0.14 76.50$\pm$1.36 86.54$\pm$0.12 89.65$\pm$0.52 92.70$\pm$0.20 92.92$\pm$0.12 96.18$\pm$0.07 77.47$\pm$0.85
GAT 64.47$\pm$0.18 76.55$\pm$1.23 86.32$\pm$0.16 90.78$\pm$0.13 93.87$\pm$0.11 93.61$\pm$0.14 96.17$\pm$0.08 76.91$\pm$0.82
APPNP 65.16$\pm$0.28 76.53$\pm$1.16 88.43$\pm$0.15 90.18$\pm$0.17 94.32$\pm$0.14 94.49$\pm$0.07 96.54$\pm$0.07 78.87$\pm$0.11
GPRGNN 67.12$\pm$0.31 77.13$\pm$1.67 89.34$\pm$0.25 89.32$\pm$0.29 94.49$\pm$0.14 95.13$\pm$0.09 96.85$\pm$0.08 78.12$\pm$0.23
GraphSAINT 67.85$\pm$0.21 -- 88.96$\pm$0.16 90.22$\pm$0.15 91.72$\pm$0.13 94.41$\pm$0.09 96.43$\pm$0.05 --
GraphSAGE -- 75.58$\pm$1.33 87.48$\pm$0.38 91.20$\pm$0.29 94.59$\pm$0.14 93.91$\pm$0.13 96.49$\pm$0.06 74.77$\pm$0.95
PPRGo 63.54$\pm$0.25 -- 87.38$\pm$0.11 88.69$\pm$0.21 93.61$\pm$0.12 92.52$\pm$0.15 95.51$\pm$0.08 78.12$\pm$0.23
GRAND+ 71.37$\pm$0.11 -- 88.64$\pm$0.09 88.74$\pm$0.11 94.75$\pm$0.12 93.92$\pm$0.08 96.47$\pm$0.04 --
GT GT 61.05$\pm$0.38 -- 88.79$\pm$0.12 91.18$\pm$0.17 94.74$\pm$0.13 94.64$\pm$0.13 97.05$\pm$0.05 --
Graphormer OOM -- OOM OOM 92.74$\pm$0.14 94.64$\pm$0.13 OOM --
SAN 59.01$\pm$0.34 -- 88.22$\pm$0.15 89.93$\pm$0.16 94.86$\pm$0.10 94.51$\pm$0.15 OOM --
GraphGPS 55.76$\pm$0.23 76.99$\pm$1.12 88.94$\pm$0.16 OOM 95.06$\pm$0.13 93.93$\pm$0.15 OOM 78.66$\pm$0.49
GOAT -- 76.89$\pm$1.19 86.87$\pm$0.24 90.96$\pm$0.90 92.96$\pm$1.48 94.21$\pm$0.38 96.24$\pm$0.24 77.00$\pm$0.77
NodeFormer -- 76.33$\pm$0.59 89.32$\pm$0.25 86.98$\pm$0.62 93.46$\pm$0.35 95.64$\pm$0.22 96.45$\pm$0.28 74.73$\pm$0.94
DIFFormer -- 76.72$\pm$0.68 89.51$\pm$0.67 91.99$\pm$0.76 95.10$\pm$0.47 94.78$\pm$0.20 96.60$\pm$0.18 73.46$\pm$0.56
NAGphormer 71.51$\pm$0.13 77.42$\pm$1.41 89.70$\pm$0.19 91.22$\pm$0.14 95.49$\pm$0.11 95.75$\pm$0.09 97.34$\pm$0.03 77.16$\pm$0.72
Exphormer 69.09$\pm$0.72 76.83$\pm$1.24 89.52$\pm$0.54 91.59$\pm$0.31 95.27$\pm$0.42 95.77$\pm$0.15 97.16$\pm$0.13 78.54$\pm$0.49
VCR-Graphormer 71.67$\pm$0.10 -- 89.77$\pm$0.15 91.75$\pm$0.15 95.53$\pm$0.14 95.37$\pm$0.04 97.34$\pm$0.04 --
GQT (ours) 71.81$\pm$0.21 77.84$\pm$0.94 90.14$\pm$0.16 93.37$\pm$0.44 95.73$\pm$0.18 96.11$\pm$0.09 97.53$\pm$0.06 80.14$\pm$0.57
Homophilic Node Classification. We use eight medium-scale homophilic datasets including: CoraFull (Bojchevski and Günnemann, 2017), CiteSeer, PubMed (Yang et al., 2016), Amazon Computers, Amazon Photos, Co-author CS, Co-author Physics (Shchur et al., 2018), and WikiCS (Mernyei and Cangea, 2020). We compare our results with eight GNNs including: GCN (Kipf and Welling, 2017), GAT, APPNP (Gasteiger et al., 2018), GPRGNN (Chien et al., 2020), GraphSAINT (Zeng et al., 2020), GraphSAGE (Hamilton et al., 2017), PPRGo (Bojchevski et al., 2020), and GTAND+ (Feng et al., 2022). We also compare against ten GTs including GT (Dwivedi and Bresson, 2020), Graphormer (Ying et al., 2021), SAN (Kreuzer et al., 2021), GraphGPS (Rampášek et al., 2022), GOAT (Kong et al., 2023), NodeFormer (Wu et al., 2022), DiffFormer (Wu et al., 2023a), NAGphormer (Chen et al., 2023), Exphormer (Shirzad et al., 2023), and VCR-Graphormer (Fu et al., 2024). The baseline performance is reported from (Wu et al., 2023b; Luo et al., 2024a). GQT outperforms the baseline GNN and GT models on 7 out of 8 benchmarks (Table 2). Notably, this achievement comes with a significant memory reduction. For example, on the Physics dataset with 34,493 nodes, we only use $256\times 6$ tokens, i.e., a 23-fold memory reduction.
Heterophilic Node Classification. We also evaluate GQT on six medium-scale heterophilic datasets: Squirrel, Chameleon (Rozemberczki et al., 2021), Questions, Roman-Empire, Amazon-Ratings, and Minesweeper (Platonov et al., 2023). We compare the performance with seven GNNs: GCN, GraphSAGE, GAT, GPRGNN, H2GCN (Zhu et al., 2020a), CPGNN (Zhu et al., 2021), and GloGNN (Li et al., 2022), and six GTs: GraphGPS, GOAT, NodeFormer, SGFormer, NAGphormer, and Exphormer. The baseline performance is reported from (Wu et al., 2023b; Luo et al., 2024b; Platonov et al., 2023; Behrouz and Hashemi, 2024). As shown in Table 3, GQT outperforms the baselines on five out of six datasets. We observe that introducing semantic edges and structural gating specifically benefits the heterophilic setting (Appendix 12.1), as they enable the Transformer to capture long-range dependencies that are not easily accessible through the original graph structure.
Large-scale Node Classification We also use four large-scale datasets: ogbn-proteins, ogbn-arxiv, ogbn-products (Hu et al., 2020a), and pokec (heterogeneous) (Leskovec and Krevl, 2014). We compare the performance against six GNN: LINKX (Lim et al., 2021), SIGN (Frasca et al., 2020), GCN, GAT, GraphSAGE, and GPRGNN; and six GTs: GraphGPS, GOAT, NodeFormer, NAGphormer, Exphormer, and SGFormer (Wu et al., 2023b). We report the baseline performance from (Wu et al., 2023b; Luo et al., 2024a). The results (Table 4) show that GQT outperforms the baseline models on all large-scale benchmarks. This achievement comes with a significant reduction in required memory. For instance, on the ogbn-products dataset with 2,449,029 nodes and 100-dimensional node features, GQT requires only 3 codebooks of size 4096, resulting in a 30-fold memory reduction.
Table 3: Mean node classification performance on heterophilic graphs over five runs.
Dataset Squirrel Chameleon Amazon-Ratings Roman-Empire Minesweeper Questions
#Nodes 5,201 2,277 22,662 24,492 10,000 48,921
#Edges 216,933 36,101 32,927 93,050 39,402 153,540
#Features 2,089 2,325 300 300 7 301
#Classes 5 5 18 5 2 2
[…]
GNN GCN 38.67$\pm$1.84 41.31$\pm$3.05 48.70$\pm$0.63 73.69$\pm$0.74 89.75$\pm$0.52 76.09$\pm$1.27
GraphSAGE 36.09$\pm$1.99 37.77$\pm$4.14 53.63$\pm$0.39 85.74$\pm$0.67 93.51$\pm$0.57 76.44$\pm$0.62
GAT 35.62$\pm$2.06 39.21$\pm$3.08 52.70$\pm$0.62 88.75$\pm$0.41 93.91$\pm$0.35 76.79$\pm$0.71
H2GCN 35.10$\pm$1.15 26.75$\pm$3.64 36.47$\pm$0.23 60.11$\pm$0.52 89.71$\pm$0.31 63.59$\pm$1.46
CPGNN 30.04$\pm$2.03 33.00$\pm$3.15 39.79$\pm$0.77 63.96$\pm$0.62 52.03$\pm$5.46 65.96$\pm$1.95
GPRGNN 38.95$\pm$1.99 39.93$\pm$3.30 44.88$\pm$0.34 64.85$\pm$0.27 86.24$\pm$0.61 55.48$\pm$0.91
GloGNN 35.11$\pm$1.24 25.90$\pm$3.58 36.89$\pm$0.14 59.63$\pm$0.69 51.08$\pm$1.23 65.74$\pm$1.19
GT GraphGPS 39.67$\pm$2.84 40.79$\pm$4.03 53.10$\pm$0.42 82.00$\pm$0.61 90.63$\pm$0.67 71.73$\pm$1.47
NodeFormer 38.52$\pm$1.57 34.73$\pm$4.14 43.86$\pm$0.35 64.49$\pm$0.73 86.71$\pm$0.88 74.27$\pm$1.46
SGFormer 41.80$\pm$2.27 44.93$\pm$3.91 48.01$\pm$0.49 79.10$\pm$0.32 90.89$\pm$0.58 72.15$\pm$1.31
NAGphormer 35.80$\pm$1.33 -- 51.26$\pm$0.72 74.34$\pm$0.77 84.19$\pm$0.66 --
Exphormer 36.04$\pm$1.45 -- 53.51$\pm$0.46 89.03$\pm$0.37 90.74$\pm$0.53 --
GQT(ours) 42.72$\pm$1.69 44.23$\pm$3.05 54.32$\pm$0.41 90.98$\pm$0.24 97.36$\pm$0.35 78.94$\pm$0.86
Table 4: Mean node classification performance on large-scale datasets over five runs.
Dataset ogbn-proteins ogbn-arxiv ogbn-products pokec
#Nodes 132,534 169,343 2,449,029 1,632,803
#Edges 39,561,252 1,166,243 61,859,140 30,622,564
#Features 128 8 100 65
#Classes 40 2 47 2
[…]
GNN GCN 72.51$\pm$0.35 71.74$\pm$0.29 75.64$\pm$0.21 75.45$\pm$0.17
GAT 72.02$\pm$0.44 71.95$\pm$0.36 79.45$\pm$0.59 72.23$\pm$0.18
GPRGNN 75.68$\pm$0.49 71.10$\pm$0.12 79.76$\pm$0.59 72.23$\pm$0.18
LINKX 71.37$\pm$0.58 66.18$\pm$0.33 71.59$\pm$0.71 82.04$\pm$0.07
GraphSAGE 77.68$\pm$0.20 71.49$\pm$0.27 78.29$\pm$0.16 75.63$\pm$0.38
SIGN -- 71.95$\pm$0.11 80.52$\pm$0.16 --
GT GraphGPS 76.83$\pm$0.26 70.97$\pm$0.41 OOM OOM
GOAT 74.18$\pm$0.37 72.41$\pm$0.40 82.00$\pm$0.43 66.37$\pm$0.94
NodeFormer 77.45$\pm$1.15 59.90$\pm$0.42 72.93$\pm$0.13 71.00$\pm$1.30
SGFormer 79.53$\pm$0.38 72.63$\pm$0.13 74.16$\pm$0.31 73.76$\pm$0.24
NAGphormer 73.61$\pm$0.33 70.13$\pm$0.55 73.55$\pm$0.21 76.59$\pm$0.25
Exphormer 74.58$\pm$0.26 72.44$\pm$0.28 OOM OOM
GQT(ours) 82.13$\pm$0.34 73.14$\pm$0.16 82.46$\pm$0.17 83.76$\pm$0.24
6.2 Ablation Study
Effect of Tokenization. We examine the performance of the tokenizer by training a linear model on the representations of the learned tokens without modulation, augmentation, or Transformer (1). As shown in Table 6, within the linear evaluation protocol, the tokenizer shows strong performance, surpassing that of GTs such as GraphGPS and NAGphormer, as well as GNNs like GAT and SIGN (Table 4). This implies that the tokenizer is capable of learning effective token representations. To further investigate the importance of the tokenizer, we exclude it and train the Transformer directly on the original node features (2). As expected, this results in significant degradation in performance, highlighting the crucial role of the tokenizer. Additionally, to study the effects of vector quantization, GraphMAE2, and DGI objectives, we train the model by excluding each component (3-5). The results suggest that the SSL objectives contribute more significantly to the performance compared to vector quantization. This is because the primary purpose of vector quantization is to compress information into discrete tokens, reducing memory requirements. Between GraphMAE2 and DGI, the former yields the highest gain. This is due to its composition of two objectives: masked reconstruction and teacher-(noisy)student distillation. Both of these objectives have been shown to outperform InfoMax objectives on downstream tasks (Hou et al., 2022; Thakoor et al., 2022).
Effect of Modulation. We also investigate the impact of codebook embeddings, positional encoding, and structural gating on the model’s performance (6-8). As shown in Table 6, introducing aggregated codebook embeddings leads to improved downstream performance because it provides the Transformer with richer representations of each token. Positional encoding, as observed in other domains, contributes moderately to downstream performance. We also note that introducing structural gating yields moderate improvements in homophilic settings, whereas the gains are significant in heterophilic benchmarks (12.1). This disparity can be attributed to the ability of structural gating to provide the Transformer with importance scores computed over the global graph structure, which is particularly beneficial in heterophilic scenarios.
Effect of Augmentation. We study the effect of semantic edges on downstream performance (9). The results suggest that augmenting the graph structure with semantic edges yields significant gains. This is because introducing semantic edges allows the Transformer to access semantic information that may not be captured by the original graph structure. Furthermore, when combined with random walks, this also enables the Transformer to attend to long-range dependencies which is particularly important in heterophilic benchmarks, where semantic relationships between nodes are more nuanced.
Table 5: Accuracy drop under adversarial attack.
Attack GR-BCD PR-BCD
PubMed ogbn-arxiv PubMed ogbn-arxiv
RQ-VAE 20.40% 14.80% 23.30% 17.20%
GQT (ours) 15.80% 10.40% 18.10% 11.30%
Robustness Analysis. To measure robustness, we use Greedy Randomized Block Coordinate Descent (GRBCD) and Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021) adversarial attacks to measure the accuracy degradation. We compare the GQT with an RQ-VAE (Lee et al., 2022). The results (Table 5) show that our tokenizer is more robust to attacks. This is because GQT is trained with multi-task self-supervised objectives while RQ-VAE is trained with reconstruction objective. More details are provided in Appendix 12.
Table 6: Ablation study on effect of proposed components on the ogbn-arxiv dataset.
Graph Tokenizer Token Modulation Augmentation Model Performance
[…]
Embeddings Encoding Gating Edges Sequence
[…]
7 Conclusion
We introduced the Graph Quantized Tokenizer (GQT) to provide standard Transformer encoders to acces discrete graph tokens that encapsulate local interactions and allow Transformers to attend to long-range dependencies within the graph structure. This allows us to seamlessly take advantage of the rapid advances in scaling Transformers. We achieved state-of-the-art performance on 20 out of 22 datasets, including large-scale and long-range homophilic and heterophilic datasets. As future directions, we plan to explore the potential of GQT in generative graph learning. Additionally, we aim to couple GQT with LLMs to provide a shared feature space across various graph datasets, paving the way for true Graph Foundational Models (GFMs) (Liu et al., 2023b; Mao et al., 2024).